# Improving Intrusion Detection with Domain-Invariant Representation Learning in Latent Space

### v2

@@ Abstract
Domain generalization focuses on leveraging knowledge from multiple related domains with ample training data and labels to enhance inference on unseen in-distribution (IN) and out-of-distribution (OOD) domains. In our study, we introduce a two-phase representation learning technique using multi-task learning. This approach aims to cultivate a latent space from features spanning multiple domains, encompassing both native and cross-domains, to amplify generalization to IN and OOD territories. Additionally, we attempt to disentangle the latent space by minimizing the mutual information between the prior and latent space, effectively de-correlating spurious feature correlations. Collectively, the joint optimization will facilitate domain-invariant feature learning. We assess the model’s efficacy across multiple cybersecurity datasets, using standard classification metrics on both unseen IN and OOD sets, and juxtapose the results with contemporary domain generalization methods.
@@ 1 Introduction
Domain generalization seeks to learn from multiple source domains and subsequently assess the model on an unseen target domain [18, 19]. A domain typically constitutes a probability distribution where training samples originate. These samples stem from a source domain, while a separate target domain is kept for testing. Deep neural networks excel when tested on data from the same distribution as their training data. Within the few-shot learning (FSL) framework, models aim to adapt to out-of-distribution (OOD) data with minimal labeled samples [10, 9, 8, 7, 6]. Meta-learning endeavors to train models across varied tasks, enabling them to grasp new tasks with few samples [11]. A pivotal challenge arises from the domain shift between training and test distributions, compromising generalization [12]. Many recent studies train with both in-distribution and OOD data, though this can infuse bias. Dataset bias emerges when a model trained on one dataset falters on test data from a distinct distribution [21]. This is commonly known as domain adaptation where information about the target domain or its classes is accessible during the training process[22]. Some approaches augment training data with perturbations mimicking domain shifts. In domain generalization, the underlying assumption is that there’s no foreknowledge of the target distribution. The challenge lies in designing a model that, once trained on single or multiple datasets, can classify data in unfamiliar domains without further training. We aspire to devise features that bridge multiple related domains without directly engaging them during the training process. Latent space models like Autoencoders function by encoding inputs into a lower-dimensional latent space and subsequently decoding them to get an estimate of the original input. A significant inspiration for our work is denoising autoencoders, which aim to reconstruct original data from noise-altered input, thereby bolstering encoder robustness. Previous research has proposed methods such as minimizing divergence between data source marginal distributions or adversarial feature alignment using maximum mean discrepancy [21]. We focus on learning a disentangled latent representation by consistently minimizing mutual information between prior and latent spaces. This action indirectly tries to maximize the mutual independence between input and latent spaces. Our model adopts a supervised autoencoder (SAE) approach [15], where both classification and reconstruction processes occur in the learned latent space simultaneously. However, our multi-domain latent space is curated with samples spanning multiple other domains(cross-domains) with the mutual invariance loss trying to remove the spurious correlation of the original high dimensional data. Recently, the work [37] has highlighted the problem of how spurious correlation can hurt generalization performance when the task correlation changes among tasks and solves the problem with multi-task causal representation learning with task-specific invariant regularization. In our case, we consider the scenario where we do not know the features with spurious correlation beforehand and therefore, directly minimize the MI content between the prior and the latent space which indirectly will aim to remove the spurious correlation in the latent space and create a disentangled latent space. This is based on the intuition that most probably the spurious correlation of the high-dimensional input space will be transferred to the latent space if there is no proper regularization to minimize it. Multi-task learning will facilitate learning representations from multiple diverse datasets and therefore jointly optimizing it with a de-correlation penalty between input and latent space will enable learning a disentangled or invariant latent space representation which will help us achieve generalization to unseen domains. In summary, our contributions can be highlighted as:
  • We introduce a two-stage representation learning approach for domain generalization in the latent space. In the first phase, we craft a multi-domain latent space by involving the data from multiple source and cross-domains. Simultaneously, we try to cultivate a disentangled or invariant latent space by applying a mutual information(MI) regularization which tries to minimize the MI between the high-dimensional prior space and the latent space. The regularization acts as a de-correlation penalty by removing spurious correlation that exists in different attack datasets in the original high-dimensional input space. These two components jointly try to enforce domain invariant learning in the latent space.
  • Our methodology jointly optimizes the classification loss, the multi-domain reconstruction loss, and the mutual invariance regularization in the latent space. This joint optimization seeks to boost classification outcomes, particularly considering both IN and OOD classes.
@@
  • We perform a wide range of experiments on cybersecurity datasets. Our results showcase overall significant improvement in generalization performance for all the source, cross-domain, and OOD classes when evaluated with standard classification metrics.
@@ 2 Related Work
1) Domain Invariant Representation Learning: This method aims to identify domain invariant representations that can be extended to unseen domains. The crux of these strategies, as seen in works such as [35], is to filter out domain-specific insights while maintaining cross-domain information. Notable studies employing autoencoders, such as [1], amalgamate multiple domains during training, augmented by data enhancement techniques, to extract domain-invariant characteristics. These features then demonstrate superior generalization to out-of-distribution data. Another study, Maximum Mean Discrepancy Adversarial Autoencoder (MMD-AAE) [3], in the context of few-shot learning, emphasizes aligning varied domain distributions to a generic prior distribution while engaging in adversarial feature learning. An innovative approach is suggested in [4], where a domain-centric masking technique is applied to learn both domain-specific and domain-invariant features. This will facilitate efficient source domain classification and sufficient generalization to target domains. In [15], a noise-enhanced supervised autoencoder is presented. This autoencoder undertakes a dual task of reconstructing inputs and classifying both the inputs and the reconstructed entities by feeding them back as input to the model. Leveraging the intra-class correlation (ICC) metric, they demonstrate the superior discriminative and generalizing capabilities of the learned feature embeddings. Moreover, authors in [26] propose domain generalization through domain-invariant representation that uniformly distributes across multiple source domains. Their approach employs moment alignment of distributions and enforces feature disentanglement via an entropy loss.
@@
3) Information Bottleneck Principle: In contrast to the aforementioned methodologies, our strategy propels direct disentanglement or decorrelation between multiple training domains. Decorrelating spurious feature correlations in order to isolate signals from background noise has been successfully used in the physics and statistics community before. An information-theoretic perspective on variance-invariance-covariance has been provided here [29] in the context of self-supervised learning which helps to achieve generalization guarantees for downstream supervised learning tasks. Adversarial learning-based domain adaptation methods are prone to negative transfer which hurts the generalization performance [28]. In our method, class-invariant features are directly learned in the latent dimension by minimizing mutual information between latent and prior space. Utilizing kernel-based metrics, we non-parametrically estimate mutual information between latent and prior space in order to achieve this.
4) Metric Learning: Metric learning aims to learn a representation function that can map higher-dimensional data to a latent embedded space. The distance in embedded space should preserve the similarity between the similar classes and the dissimilar classes are pushed away from each other. The authors [36] propose mixing target labels with training samples to improve the quality of representations or embeddings for classification purposes. In deep metric learning, we try to learn non-linear mapping from input space to low-dimensional latent space [33]. Various mix-up techniques have been recently proposed to improve the quality of representation learned by adopting some interpolation techniques between pairs of input samples and their labels [34]. On the other hand, we follow a multi-task representation learning approach that aims to learn a robust feature representation in the latent space by leveraging multiple sources and cross-domain feature information.
@@ 3 Problem Formulation
In our domain generalization problem, we consider $M$ training tasks comprising of source and cross-domain data consisting of $N$ samples, $\mathcal{S}_{train}=\{X^{i},Y^{i}|i=1,2,3...,M\}$ where, ${X}^{i},{Y}^{i}=\{\left(x^{i}_{j},y^{i}_{j}\right)\}_{j=1}^{N_{i}}$. Let $I=\{I_{1},I_{2},...,I_{k}\}$ represent the power set of these labels, where $I_{q}\subseteq\{0,1,...,K\}$. For example, if $I_{1}$ is a source domain, where, $I_{1}=\{0,1,2,3,4\}$, then, we add multiple cross-domain data $I_{2},I_{3},I_{4}$, where, $I_{2}=\{5,6,7\}$, $I_{3}=\{8,9,10\}$, and $I_{4}=\{11,12,13\}$. Let’s assume that in this context, the cardinality of each $I_{q}$, denoted as $|I_{q}|$, is constrained such that $|I_{q}|\leq 5$. The extension to multiple domains necessitates the definition of a multi-task learning objective over all the M source and cross domains which can be given as
@@ 3.1 Mutual Invariance Regularization
In order to create a domain-invariant latent space, we use a nonparametric way to estimate the mutual information between the prior and the latent space directly from the samples.
@@
3.1.1 Using Renyi’s Entropy and Joint Entropy to Calculate Mutual Information Between Inp and Latent Space
@@
In information theory, the dependence measure or the total correlation quantifies how much the feature variables either in latent space or input space, deviate from the statistical independence in each dimension and is expressed as the Kullback Leibler(KL) divergence between the joint probability distribution and the marginal distribution of the feature variables[b16]. In our case, we enforce the de-correlation between the high-dimensional input space and low-dimensional latent space representing multiple source and cross-domain data by minimizing the mutual information between the input and the latent space as a regularization technique. This will help in creating a disentangled latent space that is devoid of spurious correlations of the high-dimensional input space.
@@
For a given data matrix $X^{l}$, where $l$ represents a batch size, the matrix-based Renyi’s $\alpha$ entropy for a normalized positive semi-definite matrix ${X}$ of size $n\times n$, where $n$ represents the number of samples in each batch, can be estimated in the input space as
@@
$\begin{split}\hat{H_{2}}(X^{l})=\frac{1}{1-\alpha}\log_{2}\left[tr(X^{l})^{\alpha})\right]=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{n}\lambda_{i}(X^{l})^{\alpha}\right),\end{split}$ (2)
where $\left(X^{1}\right)_{i,j}=\mathcal{K}_{1}\left(x^{1}_{i},x^{1}_{j}\right)$, $\left(X^{2}\right)_{i,j}=\mathcal{K}_{2}\left(x^{2}_{i},x^{2}_{j}\right)$ ,…, $\left(X^{l}\right)_{i,j}=\mathcal{K}_{l}\left(x^{l}_{i},x^{l}_{j}\right)$ and $\lambda_{i}({X})$ denotes the $i^{th}$ eigenvalue of $X$, $x^{l}$ represents the $l_{th}$ batch.
@@
Similarly, Renyi’s quadratic entropy of the encoded latent space $Z$ is estimated as
@@
$\begin{split}\hat{H_{2}}(Z^{l})=\frac{1}{1-\alpha}\log_{2}\left[tr(Z^{l})^{\alpha})\right]=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{n}\lambda_{i}(Z^{l})^{\alpha}\right),\end{split}$ (3)
@@
The argument in equation (3) is called the information potential. In the above section, we use the matrix-based second-order Renyi’s entropy $\left(\alpha=2\right)$ [16] to evaluate the entropy of the latent and prior space. This can be expressed in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the Hilbert space. We use this measure to estimate the entropy of the original data space. Now, we can write the matrix-based $\alpha$ order joint entropy between latent space $Z$ and prior space $X$ as
$\hat{H_{2}}\left(X^{l},Z^{l}\right)=H_{2}\left(\frac{X^{l}\circ Z^{l}}{tr\left(X^{l}\circ Z^{l}\right)}\right),$ (4)
@@
where $\circ$ represents the Hadamard product of matrices. Based on the above definitions, we calculate the mutual information between latent and prior space with the help of the matrix-based normalized Renyi’s entropy of the latent space and the joint entropy between the prior space and latent space directly from the samples.
@@ 3.1.2 The Mutual Information Divergence
We use the non-parametric method of mutual information divergence in order to estimate the mutual information between the latent and prior space. Minimizing the mutual information indirectly results in de-correlating the feature correlation that exists in the original input space and the latent space which helps in improving the generalization performance. The mutual information during each batch of the training can be estimated as
$\begin{split}\hat{MI}(X^{l};Z^{l})&=\hat{H_{2}}\left(X^{l}\right)+\hat{H_{2}}\left(Z^{l}\right)-\hat{H_{2}}\left(X^{l},Z^{l}\right),\end{split}$ (5)
where $\hat{H_{2}}(X,Z)$ is the second-order joint entropy between latent and prior space. Minimizing this divergence as a regularization technique in the final loss objective will aid in preserving useful disentangled information in the latent space during each iteration of the training process.

### v3

# Out-of-Distribution Detection with Domain-Invariant Representations in Multi-domain Latent Space

@@ Abstract
Domain generalization focuses on leveraging the knowledge from the training data of multiple related domains to enhance inference on unseen in-distribution (IN) and out-of-distribution (OOD) domains. In our study, we introduce a multi-task representation learning technique that leverages the information of multiple related domains to improve the detection of classes from unseen domains. Our method aims to cultivate a latent space from data spanning multiple domains, encompassing both source and cross-domains, to amplify generalization to OOD domains. Additionally, we attempt to disentangle the latent space by minimizing the mutual information between the input and the latent space, effectively de-correlating spurious correlations among the samples of a specific domain. Collectively, the joint optimization will facilitate domain-invariant feature learning. We assess the model’s efficacy across multiple cybersecurity datasets, using standard classification metrics on both unseen IN and OOD sets, and validate the results with contemporary domain generalization methods.
@@ 1 Introduction
Domain generalization seeks to learn from multiple related source domains and subsequently evaluate the model on an unseen target domain [18, 19]. A domain typically constitutes a particular probability distribution from which the training samples are drawn. These samples stem from a source domain, while a separate target domain is kept for testing. Deep neural networks excel when tested on data from the same distribution as their training data. Within the few-shot learning (FSL) framework, models aim to adapt to out-of-distribution (OOD) data with minimal labeled samples [10, 8, 7, 6]. Meta-learning endeavors to train models across varied tasks, enabling them to grasp new tasks with few samples [11]. A pivotal challenge arises from the domain shift between training and test distributions, compromising generalization [12]. Many recent studies train with both in-distribution and OOD data, though this can infuse bias. In the case of domain adaptation, it is considered that the information about the target domain or its classes is accessible during the training process [22]. Some approaches augment training data with perturbations mimicking the target domain. In domain generalization, the underlying assumption is that there’s no foreknowledge of the target distribution. The challenge lies in designing a model that, once trained on single or multiple datasets, can classify data in unfamiliar domains without further training. We aspire to devise domain invariant features that bridge multiple related domains without directly engaging them during the training process. Previous research has proposed methods such as adversarial feature alignment using maximum mean discrepancy [21] to achieve domain generalization. In our case, we focus on learning a disentangled latent representation by consistently minimizing mutual information between prior and latent spaces from multiple related domains. However, our multi-domain latent space is curated with samples spanning multiple other domains(cross-domains) with the mutual invariance penalty trying to remove the spurious correlation present in individual domains. Recently, the work [37] has highlighted the problem of how spurious correlation can hurt generalization performance when the task correlation changes among the related tasks and solves the problem with multi-task causal representation learning with task-specific invariant regularization. In our case, we consider the scenario where the samples from individual domains have spurious feature correlation and therefore minimize the mutual information content between the input and the latent space which indirectly will aim to remove the spurious correlation in the individual domains. Otherwise, this is done by measuring the entropy or the uncertainty of the data on the eigenspectrum of the gram matrix of the input and latent space kernels. Multi-task learning facilitates learning representations from multiple diverse datasets and therefore jointly optimizing it with a de-correlation penalty to achieve generalization to unseen domains. In summary, our contributions can be highlighted as:
  • We consider a scenario where we know that each domain has its own spuriously correlated features which hurts the generalization performance of the model when tested on OOD domains. To address this, we propose a multi-task representation learning approach that adapts a multi-domain latent space by incorporating data from multiple sources and cross-domains. Simultaneously, we try to cultivate a disentangled or invariant latent space by applying a regularization penalty that minimizes the mutual information(MI) between the input and the latent space.
  • Our methodology jointly optimizes the classification loss, the multi-domain reconstruction loss, and the mutual invariance regularization in the latent space. The regularization, calculated in the kernel space of the input and latent space, acts as a de-correlation penalty, or otherwise, we do not want the spurious correlation of the higher dimensional features to be projected to the latent space.
@@
  • We show that cross-domain data when added in a principled way, can improve generalization performance on the IN and OOD classes. We perform a wide range of experiments on cybersecurity datasets to showcase overall improvement in generalization performance across IN, OOD, and cross-domains classes.
@@ 2 Related Work
1) Domain Invariant Representation Learning: This method aims to identify domain invariant representations that can be extended to unseen domains. The crux of these strategies, as seen in works such as [35], is to filter out domain-specific insights while maintaining cross-domain information. Notable studies employing autoencoders, such as [1], amalgamate multiple domains during training, augmented by data enhancement techniques, to extract domain-invariant characteristics. These features then demonstrate superior generalization to out-of-distribution data. Another study, Maximum Mean Discrepancy Adversarial Autoencoder (MMD-AAE) [3], in the context of few-shot learning, emphasizes aligning varied domain distributions to a generic prior distribution while engaging in adversarial feature learning. An innovative approach is suggested in [4], where a domain-centric masking technique is applied to learn both domain-specific and domain-invariant features. This will facilitate efficient source domain classification and sufficient generalization to target domains. In [15], a noise-enhanced supervised autoencoder is presented. This autoencoder undertakes a dual task of reconstructing inputs and classifying both the inputs and the reconstructed entities by feeding them back as input to the model. Leveraging the intra-class correlation (ICC) metric, they demonstrate the superior discriminative and generalizing capabilities of the learned feature embeddings. Moreover, authors in [26] propose domain generalization through domain-invariant representation that uniformly distributes across multiple source domains. Their approach employs moment alignment of distributions and enforces feature disentanglement via an entropy loss. Recently, the DIFEX [9] paper utilizes a knowledge distillation framework to capture the high-level Fourier phase as internally-invariant characteristics, while simultaneously learning cross-domain correlation alignment to extract mutually-invariant features.
@@
3) Information Bottleneck Principle and Metric Learning: In contrast to the aforementioned methodologies, our strategy propels direct disentanglement or decorrelation between multiple training domains. An information-theoretic perspective on variance-invariance-covariance has been provided here [29] in the context of self-supervised learning which helps to achieve generalization guarantees for downstream supervised learning tasks. Adversarial learning-based domain adaptation methods are prone to negative transfer which hurts the generalization performance [28]. In our method, class-invariant features are indirectly learned in the latent dimension by minimizing mutual information between latent and prior space. Metric learning aims to learn a representation function that can map higher-dimensional data to a latent embedded space. The distance in embedded space should preserve the similarity between the similar classes and the dissimilar classes are pushed away from each other. The authors [36] propose mixing target labels with training samples to improve the quality of representations or embeddings for classification purposes. In deep metric learning, we try to learn non-linear mapping from input space to low-dimensional latent space [33]. Various mix-up techniques have been recently proposed to improve the quality of representation learned by adopting some interpolation techniques between pairs of input samples and their labels [34]. On the other hand, we follow a multi-task representation learning approach that aims to learn a robust feature representation in the latent space by leveraging multiple sources and cross-domain feature information.
@@ 3 Problem Formulation
In our domain generalization problem, we consider $M$ training tasks comprising of source and cross-domain data consisting of $N$ samples, $\mathcal{S}_{train}=\{X^{i},Y^{i}|i=1,2,3...,M\}$ where, ${X}^{i},{Y}^{i}=\{\left(x^{i}_{j},y^{i}_{j}\right)\}_{j=1}^{N_{i}}$. Let $I=\{I_{1},I_{2},...,I_{k}\}$ represent the power set of these labels, where $I_{q}\subseteq\{0,1,...,K\}$. For example, if $I_{1}$ is a source domain, where, $I_{1}=\{0,1,2,3,4\}$, then, we add multiple cross-domain data $I_{2},I_{3},I_{4}$, where, $I_{2}=\{5,6,7\}$, $I_{3}=\{8,9,10\}$, and $I_{4}=\{11,12,13\}$. The extension to multiple domains necessitates the definition of a multi-task learning objective over all the M source and cross domains which can be given as
@@ 3.1 Mutual Invariance Regularization
In order to learn a domain-invariant latent space, we use a kernel-based mutual information minimization between the prior and the latent space as a regularization technique.
@@
3.1.1 Kernel-based Renyi’s Entropy and Joint Entropy to measure Mutual Information between Input and Latent Space
@@
In information theory, the dependence measure or the total correlation between the feature variables is measured as the statistical independence in each dimension and is expressed as the Kullback Leibler(KL) divergence between the joint probability distribution and the marginal distribution of the features [16]. In our case, we enforce the de-correlation between the input and the latent kernel space comprising multiple source and cross-domains by minimizing the mutual information between them as a regularization penalty. This will try to introduce a disentangled latent space that is devoid of the spurious correlations in the individual domains in their high-dimensional input space.
@@
The matrix-based Renyi’s second-order entropy [16] of a normalized positive definite(NPD) matrix $\mathcal{K}_{x}$, of size $l\times l$ in the input space, where $l$ represents a batch size, can be estimated as
@@
$\begin{split}\hat{H_{2}}(\mathcal{K}_{x})=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{l}\lambda_{i}(\mathcal{K}_{x})^{\alpha}\right),\end{split}$ (2)
where the Gram matrix $\mathcal{K}_{x}$ is obtained by evaluating the positive definite (PSD) kernel on all $l$ pairs of samples in a batch of training, that is, $\left(\mathcal{K}_{x}\right)_{i,j}=\mathcal{K}_{x}\left(x_{i},x_{j}\right)$ and $\lambda_{i}({X})$ denotes the $i^{th}$ eigenvalue of the input kernel matrix $\mathcal{K}_{x}$ of the $l_{th}$ batch, $\mathcal{K}_{i,j}=\frac{1}{l}\frac{\mathcal{K}_{i,j}}{\sqrt{\mathcal{K}_{i,i}\mathcal{K}_{j,j}}}$ Here, $\alpha=2$ considering Renyi’s second-order entropy.
@@
Similarly, Renyi’s quadratic entropy of the latent space kernel $\mathcal{K}_{Z}$ of size $l\times l$ is estimated as
@@
$\begin{split}\hat{H_{2}}(\mathcal{K}_{z})=\frac{1}{1-\alpha}\log_{2}\left(\sum_{i=1}^{l}\lambda_{i}(\mathcal{K}_{z})^{\alpha}\right),\end{split}$ (3)
@@
The argument in equation (3) is called the information potential. In the above section, we use the matrix-based second-order Renyi’s entropy $\left(\alpha=2\right)$ [16] to evaluate the entropy or the uncertainty of the latent and the input space in terms of the normalized eigenspectrum of the Hermitian matrix of the projected data in the Hilbert space. Now, we can estimate the matrix-based second-order joint entropy between latent space kernel $Z$ and the input space kernel $X$ as
$\hat{H_{2}}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right)=H_{2}\left(\frac{\mathcal{K}_{x}\circ\mathcal{K}_{Z}}{tr\left(\mathcal{K}_{X}\circ\mathcal{K}_{z}\right)}\right),$ (4)
@@
where $\circ$ represents the Hadamard product. Based on the above definitions, we calculate the joint entropy of the latent and the input space with the help of the matrix-based normalized Renyi’s entropy of the latent space and the input space kernels. The joint entropy is used to derive the mutual information between the input and the latent space.
@@ 3.1.2 The Mutual Information Divergence
We use the matrix-based mutual information divergence to estimate the mutual information between the latent and input space kernels. Minimizing the mutual information indirectly results in de-correlating the feature correlation that exists in the original input space which helps in improving the generalization performance. The mutual information during each batch of the training can be estimated as
$\begin{split}\hat{MI}(\mathcal{K}_{X};\mathcal{K}_{Z})&=\hat{H_{2}}\left(\mathcal{K}_{X}\right)+\hat{H_{2}}\left(\mathcal{K}_{Z}\right)-\hat{H_{2}}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right),\end{split}$ (5)
where $\hat{H_{2}}\left(\mathcal{K}_{X},\mathcal{K}_{Z}\right),$ is the second-order joint entropy between the latent and the input kernel space. Minimizing this divergence as a regularization penalty in the final loss objective will aid in preserving useful disentangled information in the latent space during each iteration of the training process.